OpenStudy (anonymous):

Without drawing the graph of the equation, determine how many points the parabola has in common with the x-axis and whether its vertex lies above, on, or below the x-axis. y = –2x2 + x + 3

OpenStudy (anonymous):

set y=0? the find derivative of y

OpenStudy (anonymous):

common with x-axis, x=3/2, x=-1.........dy/dx=-4x+1=0, solve for x=1/4, plug 1/4 back to equation y, and if y is positive, above/ negative below, 0 on

OpenStudy (anonymous):

A. 1 point in common; vertex on x-axis B. 2 points in common; vertex below x-axis C. 2 points in common; vertex above x-axis D. no points in common; vertex below x-axis

OpenStudy (anonymous):

C